OpenStudy (anonymous):

Can someone help me solve these word problems? The sum of two numbers is 80 and their difference is 22.

OpenStudy (31356):

@BigBoyLP

OpenStudy (31356):

Hope that helps! Medal?

OpenStudy (timaashorty):

Sum The sum of two numbers is 80: \[x + y = 80 \] their difference is 22. x-y = 22. What two numbers add 80 and they also subract into 22?

OpenStudy (acxbox22):

solve the system of equations: x+y=80 x-y=22 solve it and you don't know how, ask

OpenStudy (anonymous):

Idk how to do any of it

OpenStudy (31356):

Thanks for the medal! :D

OpenStudy (acxbox22):

okay then just try to think what two numbers add up to 80 and when subtracted equal 22

OpenStudy (anonymous):

lol No problem

OpenStudy (timaashorty):

\[x = 51 ; y= 29\] \[51 + 29 = 80\] \[51-29 = 22\]

OpenStudy (timaashorty):

do you understand it now?

OpenStudy (anonymous):

Yes i do thank you

OpenStudy (timaashorty):

you're welcome